Consider a particle confined in a 1D box of length L. The energy spacing between any two consequtive energy level is propotional to
1
Expert's answer
2019-05-10T11:25:25-0400

Energy spectrum of particle in a 1D box:


"E_n = \frac{n^2 \pi^2 \hbar^2}{2mL^2}"

Energy spacing:


"E_{n+1} - E_n = \frac{\pi^2 \hbar^2}{2mL^2}(2n+1)"

Therefore, at large n, energy spacing between any two consequtive energy level is propotional to n